# Integral representations for the joint survival functions of the cumulated components of multinomial random vectors

### v1

@@ Abstract
This paper presents a multivariate normal integral expression for the joint survival function of the cumulated components of any multinomial random vector. This result can be viewed as a multivariate analog of Equation (7) from Carter and Pollard [2], who improved Tusnády’s inequality. Our findings are based on a crucial relationship between the joint survival function of the cumulated components of any multinomial random vector and the joint cumulative distribution function of a corresponding Dirichlet distribution. We offer two distinct proofs: the first expands the logarithm of the Dirichlet density, while the second employs Laplace’s method applied to the Dirichlet integral.
@@
Dirichlet distribution, Gaussian approximation, Laplace’s method, multinomial distribution, multivariate normal distribution, quantile coupling
MSC:
[2020]Primary: 62E17 Secondary: 62H10, 62H12, 62E20
@@ 1 Introduction
$p_{n}(\boldsymbol{k})=\frac{n!}{(n-\|\boldsymbol{k}\|_{1})!\prod_{i=1}^{d}k_{i}!}\,p_{d+1}^{n-\|k\|_{1}}\prod_{i=1}^{d}p_{i}^{k_{i}},$
where $p_{d+1}=1-\|\boldsymbol{p}\|_{1}\in(0,1)$ and $n\in\mathbb{N}$. The covariance matrix of the multinomial distribution is well-known to be $n\,\Sigma_{\boldsymbol{p}}$, where $\Sigma_{\boldsymbol{p}}=\text{diag}(\boldsymbol{p})-\boldsymbol{p}\boldsymbol{p}^{\top}$; see, e.g., [7, p. 377]. From Theorem 1 of Tanabe and Sagae [9], it is also well known that $\det(\Sigma_{\boldsymbol{p}})=p_{1}\dots p_{d}\,p_{d+1}$. The centered multivariate normal density with the same covariance structure as the above multinomial distribution is defined, for all $\boldsymbol{x}\in\mathbb{R}^{d}$, by
@@
Let $[d]=\{1,\ldots,d\}$. Let $I_{1}=(0,p_{1}]$ and $I_{j}=(p_{1}+\dots+p_{j-1},p_{1}+\dots+p_{j}]$ for all $j\in[d]\backslash\{1\}$. If $U_{1},\dots,U_{n}\smash{\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}}\mathrm{Uniform}(0,1)$, then
$\boldsymbol{X}=(X_{1},\ldots,X_{d})\equiv\Big{(}\sum_{i=1}^{n}\mathds{1}_{\{U_{i}\in I_{1}\}},\ldots,\sum_{i=1}^{n}\mathds{1}_{\{U_{i}\in I_{d}\}}\Big{)}\sim\mathrm{Multinomial}\hskip 0.56905pt(n,\boldsymbol{p}).$
If $U_{(1)}\leq\dots\leq U_{(n)}$ denote the order statistics, then
@@
$\qquad=\mathsf{P}(U_{(k_{1}+\dots+k_{i})}\leq p_{1}+\dots+p_{i},~{}\forall i\in[d])$
$\qquad=\int_{0}^{p_{1}}\int_{u_{1}}^{p_{1}+p_{2}}\dots\int_{u_{d-1}}^{p_{1}+\dots+p_{d}}n!\prod_{i=1}^{d+1}\frac{(u_{i}-u_{i-1})^{k_{i}-k_{i-1}-1}}{(k_{i}-k_{i-1}-1)!}\hskip 1.0pt\mathrm{d}\hskip 0.75ptu_{1}\hskip 1.0pt\mathrm{d}\hskip 0.75ptu_{2}\dots\hskip 1.0pt\mathrm{d}\hskip 0.75ptu_{d},$
where one defines $u_{0}=0$, $u_{d+1}=1$, $k_{0}=0$, $k_{d+1}=n+1$ in the multidimensional integral. After applying the change of variables $s_{i}=u_{i}-u_{i-1}$ and $j_{i}=k_{i}-k_{i-1}$, for all $i\in[d]$, together with $j_{d+1}=k_{d+1}-k_{d}=(n+1)-\sum_{i=1}^{d}j_{i}$, the above can be rewritten as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=\int_{0}^{p_{1}}\int_{0}^{(p_{1}-s_{1})+p_{2}}\dots\int_{0}^{\sum_{k=1}^{d-1}(p_{k}-s_{k})+p_{d}}n!\prod_{i=1}^{d+1}\frac{s_{i}^{j_{i}-1}}{(j_{i}-1)!}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
This neat equation provides a direct relationship between the joint survival function of the cumulated components of any multinomial random vector and the joint cumulative distribution function of a corresponding Dirichlet distribution.
In turn, letting $N=n-d$, $J_{i}=j_{i}-1$ for all $i\in[d+1]$, and defining the region
$\mathcal{R}_{d}=\left\{\boldsymbol{s}\in\mathcal{S}_{d}:(s_{1},s_{2},\dots,s_{i})\in\left(\sum_{k=1}^{i}p_{k}\right)\mathcal{S}_{i}~{}~{}\forall i\in[d]\right\},$ (1)
@@
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])=\int_{\mathcal{R}_{d}}\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\prod_{i=1}^{d+1}s_{i}^{J_{i}}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$ (2)
2 Quantile coupling
For any integer $m\in\mathbb{N}$, let $\lambda_{m}$ denote the error term in Stirling’s approximation for $\ln(m!)$. It is well known that, for all $m\in\mathbb{N}$,
@@ 2 Quantile coupling
Together with the notation introduced in Section 1, define, for all $i\in[d+1]$,
@@
and also
$\Lambda_{N}$ $=\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{i},\qquad\Delta_{N}=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}+\Lambda_{N}-\frac{1}{2}\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon}),$ (4)
@@
Theorem 1 below expresses the joint survival function of the cumulated components of any multinomial random vector in terms of a multivariate normal integral. The result can be seen as a multivariate analog of Eq. (7) of Carter and Pollard [2], who improved on Tusnády’s inequality. For an overview of the literature on Tusnády’s inequality and the most recent improvement in the bulk, see Massart [4] and Ouimet [6], respectively.
Theorem 1.
For all $\boldsymbol{k}\in\mathbb{N}_{0}^{d}\cap n\mathcal{S}_{d}$, one has
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
Useful alternative expressions for $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ and $\gamma^{\star}(\boldsymbol{s})$ can be found in (10) and (11), respectively.
@@
Theorem 1 also complements the local limit theorem for the multinomial distribution developed independently by Siotani and Fujikoshi [8] and Ouimet [5], and the related non-asymptotic approximations for Pearson’s chi square statistic by Bax and Ouimet [1].
@@ 3 Proofs
Note that $\sum_{i=1}^{d+1}J_{i}=N$. By taking the logarithm of the integrand in (2) and applying Stirling’s formula (3), one obtains
@@
$\quad=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}-\frac{1}{2}\ln\left\{\left(\frac{2\pi}{N}\right)^{d}\,\prod_{i=1}^{d+1}(J_{i}/N)\right\}+\sum_{i=1}^{d+1}J_{i}\ln\left(\frac{s_{i}}{J_{i}/N}\right)+\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{i}$
@@
Using the Taylor expansion
@@
and the fact that $\widetilde{\varepsilon}_{d+1}=-\sum_{i=1}^{d}\widetilde{\varepsilon}_{i}$, one can write
$\sum_{i=1}^{d+1}(J_{i}/N)\ln\left(\frac{p_{i}}{J_{i}/N}\right)=-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\widetilde{\gamma}(\boldsymbol{\varepsilon}),$ (8)
where the quantity $\widetilde{\gamma}(\boldsymbol{\varepsilon})$, as defined in (5), can be expanded as follows:
@@
$=\frac{1}{2}\sum_{i,j=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\left\{\frac{1}{p_{i}}\mathds{1}(i=j)+\frac{1}{p_{d+1}}\right\}-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}$
$\quad-\frac{1}{6}\sum_{i,j,k=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\left\{\frac{1}{p_{i}^{2}}\mathds{1}(i=j=k)-\frac{1}{p_{d+1}^{2}}\right\}$
$\quad+\frac{1}{12}\sum_{i,j,k,\ell=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\widetilde{\varepsilon}_{\ell}\left\{\frac{1}{p_{i}^{3}}\mathds{1}(i=j=k=\ell)+\frac{1}{p_{d+1}^{3}}\right\}+\mathcal{O}_{d,\boldsymbol{p}}\big{(}\|\boldsymbol{\varepsilon}\|_{1}^{5}\big{)}.$
Moreover, by [9, Eq. 21], it is known that $(\Sigma_{\boldsymbol{p}}^{-1})_{ij}=p_{i}^{-1}\mathds{1}_{\{i=j\}}+p_{d+1}^{-1}$ for all $i,j\in[d]$, so
$\frac{1}{2}\sum_{i,j=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\left\{\frac{1}{p_{i}}\mathds{1}(i=j)+\frac{1}{p_{d+1}}\right\}-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}=0.$ (9)
Therefore, using the last two equations, one can rewrite $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ as follows:
$\widetilde{\gamma}(\boldsymbol{\varepsilon})$ $=-\frac{1}{6}\sum_{i,j,k=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\left\{\frac{1}{p_{i}^{2}}\mathds{1}(i=j=k)-\frac{1}{p_{d+1}^{2}}\right\}$ (10)
$\quad+\frac{1}{12}\sum_{i,j,k,\ell=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\widetilde{\varepsilon}_{\ell}\left\{\frac{1}{p_{i}^{3}}\mathds{1}(i=j=k=\ell)+\frac{1}{p_{d+1}^{3}}\right\}+\mathcal{O}_{d,\boldsymbol{p}}\big{(}\|\boldsymbol{\varepsilon}\|_{1}^{5}\big{)}.$
Now, consider $\boldsymbol{s}=(s_{1},\ldots,s_{d})\in\mathcal{R}_{d}$ as defined in (1), and let
$\delta_{i}(s)=\frac{s_{i}-p_{i}}{p_{i}},~{}~{}i\in[d],\qquad\delta_{d+1}(s)=\frac{s_{d+1}-p_{d+1}}{p_{d+1}}=-\frac{\sum_{i=1}^{d}(s_{i}-p_{i})}{p_{d+1}}.$
In particular, note that this definition yields $\sum_{i=1}^{d+1}p_{i}\delta_{i}(s_{i})=0$. Upon applying a second-order version of the Taylor expansion
$\ln(1+x)=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}+\mathcal{O}_{\eta}(x^{4}),\quad|x|\leq\eta<1,$
with Lagrange remainder, there exists a vector $\boldsymbol{s}^{\star}=(s_{1}^{\star},\ldots,s_{d}^{\star})$ on the line segment joining $\boldsymbol{s}$ and $\boldsymbol{p}=(p_{1},\ldots,p_{d})$ such that

### v2

# Normal integral representation for the joint survival function of the cumulative sums of the components of multinomial random vectors

@@ Abstract
This paper presents a multivariate normal integral representation for the joint survival function of the cumulative sums of the components of any multinomial random vector at interior lattice points. This result can be viewed as a multivariate analog of Equation (7) in Carter and Pollard [2], whose proof starts from the beta integral representation of binomial survival probabilities and uses Laplace’s method to improve Tusnády’s inequality. Our findings are based on a crucial relationship between the joint survival function of the cumulative sums of the components of any multinomial random vector and a Dirichlet probability over a corresponding cumulative-sum region. The main motivation is that such an explicit formula may eventually help streamline the conditional quantile-transformation arguments used in the multivariate KMT approximation of Einmahl [4], a connection left for future work. We provide numerical checks of the identity for $d=2,3,4,5$.
@@
Dirichlet distribution, Gaussian integral representation, Laplace’s method, multinomial distribution, multivariate normal distribution, normal integral representation
2020 MSC:
Primary: 62E17 Secondary: 62H10, 62H12, 62E20
@@ 1 Introduction
One-dimensional quantile couplings between discrete distributions and Gaussian distributions are a central tool in probability and statistics. A classical example appears in the Komlós–Major–Tusnády (KMT) approximation [see, e.g., 6, 7, 3, 1, 8], where the empirical distribution function is coupled with a Brownian bridge. As explained by Carter and Pollard [2], a key ingredient in that construction is the quantile coupling between a $\mathrm{Binomial}\hskip 0.56905pt(n,1/2)$ random variable and a normal random variable with the same mean and variance. More precisely, if $X\sim\mathrm{Binomial}\hskip 0.56905pt(n,1/2)$ and $Y\sim\mathcal{N}(n/2,n/4)$, then one defines cutpoints $\infty=\beta_{0}<\beta_{1}<\cdots<\beta_{n}<\beta_{n+1}=\infty$ by
$\mathsf{P}(X\geq k)=\mathsf{P}(Y>\beta_{k}),\qquad k=0,1,\ldots,n.$
Sharp control of the difference between the binomial quantiles and their Gaussian counterparts is the content of Tusnády-type inequalities. The original inequality of Tusnády [16] and its later refinements [9, 2, 11] show that the normal approximation to these quantiles is much more accurate than what follows from a crude central limit theorem.
The starting point of Carter and Pollard [2] is not a local approximation of individual binomial probabilities, but rather an exact integral representation for the whole binomial tail:
$\mathsf{P}(X\geq k)=\frac{n!}{(k-1)!(n-k)!}\int_{0}^{1/2}t^{k-1}(1-t)^{n-k}\thinspace\mathrm{d}\thinspace t,\qquad 1\leq k\leq n.$ (1)
This beta integral representation converts the discrete tail probability into a continuous integral. It is the one-dimensional prototype for the Dirichlet integral representation developed in the present paper.
Now, let us see how the Gaussian structure enters. For $n/2<k<n$, let K=k1K=k-1, N=n1N=n-1, and $\varepsilon=(2K-N)/N$. Define
$2H(t)=(1+\varepsilon)\ln t+(1-\varepsilon)\ln(1-t),\qquad h(s)=H\left(\frac{1-s}{2}\right)-H(1/2),$
and let $\lambda_{m}$ denote the remainder term in Stirling’s formula, as in (5) below. With
$\gamma(\varepsilon)=\frac{(1+\varepsilon)\ln(1+\varepsilon)+(1-\varepsilon)\ln(1-\varepsilon)-\varepsilon^{2}}{2\varepsilon^{4}},\qquad\varepsilon\neq 0,$
where $\gamma(0)=1/12$ by continuous extension, and
$a_{N}(\varepsilon)=\ln(1+N^{-1})+\lambda_{N}-\lambda_{K}-\lambda_{N-K}-\frac{1}{2}\ln(1-\varepsilon^{2})-N\varepsilon^{4}\gamma(\varepsilon),$
Equation (7) of Carter and Pollard [2] can be written as
$\mathsf{P}(X\geq k)=e^{a_{N}(\varepsilon)}\sqrt{\frac{N}{2\pi}}\int_{0}^{1}\exp\left\{Nh(s)-\frac{N\varepsilon^{2}}{2}\right\}\thinspace\mathrm{d}\thinspace s.$
Near its maximum at $s=0$, the function $h$ starts as $h(s)=-\varepsilon s-s^{2}/2$ plus higher order terms. Thus the leading part of the integral is governed by the Gaussian expression
$\sqrt{\frac{N}{2\pi}}\int_{0}^{\infty}\exp\left\{-\frac{N(s+\varepsilon)^{2}}{2}\right\}\thinspace\mathrm{d}\thinspace s=\overline{\Phi}(\varepsilon\sqrt{N}),$
where $\overline{\Phi}(x)=\mathsf{P}(Z\geq x)$ for $Z\sim\mathcal{N}(0,1)$. Carter and Pollard [2] then use this representation, together with careful Taylor bounds and inequalities for ratios of normal tail probabilities, to sharpen Tusnády’s inequality.
The goal of the present paper is to develop the corresponding integral representation in the multinomial setting, with the hope that such an explicit formula may eventually help simplify the difficult conditional quantile-transformation arguments used in the multivariate KMT approximation of Einmahl [4]; see also Zaitsev [17]. If $\boldsymbol{X}=(X_{1},\ldots,X_{d})\sim\mathrm{Multinomial}\hskip 0.56905pt(n,\boldsymbol{p})$, the natural multivariate analog of a binomial tail is the joint survival function of the cumulative sums of the components
$\mathsf{P}(X_{1}+\cdots+X_{i}\geq k_{1}+\cdots+k_{i},~\forall i\in[d]),$
where $d]\equiv\{1,\ldots,d\}$. These probabilities arise naturally when multinomial counts are constructed from a partition of the unit interval. In that construction, the event above can be expressed in terms of selected order statistics of independent uniform random variables; see (2) below. The joint density of those order statistics yields a Dirichlet-type integral over a region $\mathcal{R}_{d}$ inside the simplex; see (4) below. When $d=1$ and $p_{1}=1/2$, this identity reduces exactly to the beta integral representation for binomial tails in (1).
The main result of the paper (Theorem 1) is an exact multivariate normal integral representation for this joint survival function at interior lattice points. This identity is not an asymptotic approximation. It is an exact rewriting of a Dirichlet integral, obtained by applying Stirling’s formula and completing the square in the natural covariance structure of the multinomial distribution. In this sense, it is a multivariate analog of the normal integral representation used by Carter and Pollard [2] in the binomial case to improve Tusnády’s inequality.
The rest of the paper is organized as follows. Section 2 introduces the necessary notation and derives the Dirichlet representation of the joint survival function of the cumulative sums of multinomial components. Section 3 states the multivariate normal integral representation. Section 4 gives two proofs: the first expands the logarithm of the Dirichlet density directly, while the second uses a Laplace-type decomposition of the Dirichlet integral. Section 5 provides numerical checks of the claimed identity.
2 Definitions and notation
@@
$p_{n}(\boldsymbol{k})=\frac{n!}{(n-\|\boldsymbol{k}\|_{1})!\prod_{i=1}^{d}k_{i}!}\,p_{d+1}^{n-\|\boldsymbol{k}\|_{1}}\prod_{i=1}^{d}p_{i}^{k_{i}},$
where $p_{d+1}=1-\|\boldsymbol{p}\|_{1}\in(0,1)$ and $n\in\mathbb{N}$. Throughout, when $\boldsymbol{s}\in\mathcal{S}_{d}$, set $s_{d+1}=1-\|\boldsymbol{s}\|_{1}$. The covariance matrix of the multinomial distribution is well-known to be $n\,\Sigma_{\boldsymbol{p}}$, where $\Sigma_{\boldsymbol{p}}=\text{diag}(\boldsymbol{p})-\boldsymbol{p}\boldsymbol{p}^{\top}$; see, e.g., [13, p. 377]. From Theorem 1 of Tanabe and Sagae [15], it is also well known that $\det(\Sigma_{\boldsymbol{p}})=p_{1}\dots p_{d}\,p_{d+1}$. The centered multivariate normal density with covariance matrix $\Sigma_{\boldsymbol{p}}$ (the per-trial multinomial covariance matrix) is defined, for all $\boldsymbol{x}\in\mathbb{R}^{d}$, by
@@
[…]
$\boldsymbol{X}=(X_{1},\ldots,X_{d})\equiv\Big(\sum_{i=1}^{n}\mathds{1}\{U_{i}\in I_{1}\},\ldots,\sum_{i=1}^{n}\mathds{1}\{U_{i}\in I_{d}\}\Big)\sim\mathrm{Multinomial}\hskip 0.56905pt(n,\boldsymbol{p}).$
For any $\boldsymbol{k}\in\mathbb{N}^{d}\cap n\mathcal{S}_{d}$, set $K_{0}=0$, $K_{i}=k_{1}+\dots+k_{i}$ for all $i\in[d]$, and $K_{d+1}=n+1$. If $U_{(1)}\leq\dots\leq U_{(n)}$ denote the order statistics, then
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~\forall i\in[d])$ (2)
$\qquad=\mathsf{P}(U_{(K_{i})}\leq p_{1}+\dots+p_{i},~\forall i\in[d])$
$\qquad=\int_{0}^{p_{1}}\int_{u_{1}}^{p_{1}+p_{2}}\dots\int_{u_{d-1}}^{p_{1}+\dots+p_{d}}n!\prod_{i=1}^{d+1}\frac{(u_{i}-u_{i-1})^{K_{i}-K_{i-1}-1}}{(K_{i}-K_{i-1}-1)!}\thinspace\mathrm{d}\thinspace u_{1}\thinspace\mathrm{d}\thinspace u_{2}\dots\thinspace\mathrm{d}\thinspace u_{d},$
where one defines $u_{0}=0$ and $u_{d+1}=1$ in the multidimensional integral. After applying the change of variables $s_{i}=u_{i}-u_{i-1}$ for all $i\in[d]$, setting $s_{d+1}=1-\|\boldsymbol{s}\|_{1}$, and writing $k_{d+1}=K_{d+1}-K_{d}=(n+1)-\|\boldsymbol{k}\|_{1}$, the above can be rewritten as
@@
$\qquad=\int_{0}^{p_{1}}\int_{0}^{(p_{1}-s_{1})+p_{2}}\dots\int_{0}^{\sum_{k=1}^{d-1}(p_{k}-s_{k})+p_{d}}n!\prod_{i=1}^{d+1}\frac{s_{i}^{k_{i}-1}}{(k_{i}-1)!}\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$
This identity provides a direct relationship between the joint survival function of the cumulative sums of the components of any multinomial random vector and a Dirichlet probability over the corresponding cumulative-sum region.
In turn, letting N=ndN=n-d, $J_{i}=k_{i}-1$ for all $i\in[d+1]$, so that $J_{d+1}=n-\|\boldsymbol{k}\|_{1}$, writing $\boldsymbol{J}=(J_{1},\ldots,J_{d})^{\top}$, and defining the regions
$\mathcal{R}_{d}=\left\{\boldsymbol{s}\in\mathcal{S}_{d}:(s_{1},s_{2},\dots,s_{i})\in\left(\sum_{k=1}^{i}p_{k}\right)\mathcal{S}_{i}~~\forall i\in[d]\right\},\qquad\mathcal{R}_{d}^{\circ}=\mathcal{R}_{d}\cap\mathrm{Int}(\mathcal{S}_{d}),$ (3)
@@
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~\forall i\in[d])=\int_{\mathcal{R}_{d}}\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\prod_{i=1}^{d+1}s_{i}^{J_{i}}\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$ (4)
[…]
3 Normal integral representation
For any integer $m\in\mathbb{N}$, let $\lambda_{m}$ denote the error term in Stirling’s approximation for $\ln(m!)$:
@@ 3 Normal integral representation
For the normal integral representation, assume from now on that
$\boldsymbol{k}\in\mathcal{K}_{n,d}=\left\{\boldsymbol{k}\in\mathbb{N}^{d}\cap n\mathcal{S}_{d}:k_{i}\geq 2~~\forall i\in[d],~\|\boldsymbol{k}\|_{1}\leq n-1\right\}.$
This condition is equivalent to $J_{i}\in\mathbb{N}$ for all $i\in[d+1]$, and it ensures that $N\in\mathbb{N}$ and that all logarithms below are finite. Together with the notation introduced in Section 1, define, for all $i\in[d+1]$,
@@
and write
$\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{d+1}),\qquad\widetilde{\boldsymbol{\varepsilon}}=(\widetilde{\varepsilon}_{1},\ldots,\widetilde{\varepsilon}_{d})^{\top}.$
Also, set
$\Lambda_{N}=\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{J_{i}},$
and
$\Delta_{N}$ $=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}+\Lambda_{N}-\frac{1}{2}\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon}),$ (6)
@@
Theorem 1 below expresses the joint survival function of the cumulative sums of the components of any multinomial random vector at interior lattice points $\boldsymbol{k}\in\mathcal{K}_{n,d}$ in terms of a multivariate normal integral. The result can be seen as a multivariate analog of Eq. (7) of Carter and Pollard [2], who improved on Tusnády’s inequality. For an overview of the literature on Tusnády’s inequality and the most recent improvement in the bulk, see Massart [9] and Ouimet [11], respectively.
Theorem 1 (Normal integral representation).
For all $\boldsymbol{k}\in\mathcal{K}_{n,d}$, one has
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~\forall i\in[d])=e^{\Delta_{N}}\int_{\mathcal{R}_{d}^{\circ}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\thinspace\mathrm{d}\thinspace\boldsymbol{s}.$
When $\max_{i\in[d+1]}|\varepsilon_{i}|\leq\eta<1$, a useful expansion for $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ can be found in (12). An exact alternative expression for $\gamma^{\star}(\boldsymbol{s})$ can be found in (13).
@@
The restriction $\boldsymbol{k}\in\mathcal{K}_{n,d}$ is used only for the normal representation above. The Dirichlet identity (4) remains valid when some $J_{i}=0$, with the convention $0!=1$, but those boundary cases are not covered by Theorem 1 because the logarithms in (6), (7), and (8) would not all be finite.
Remark 2.
Theorem 1 also complements the local limit theorem for the multinomial distribution developed independently by Siotani and Fujikoshi [14] and Ouimet [10].
@@ 4 Proofs
Under the assumption $\boldsymbol{k}\in\mathcal{K}_{n,d}$, one has $J_{i}\in\mathbb{N}$ for all $i\in[d+1]$ and $\sum_{i=1}^{d+1}J_{i}=N$. For $\boldsymbol{s}\in\mathcal{R}_{d}^{\circ}$, by taking the logarithm of the integrand in (4) and applying Stirling’s formula (5), one obtains
@@
$\quad=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}-\frac{1}{2}\ln\left\{\left(\frac{2\pi}{N}\right)^{d}\,\prod_{i=1}^{d+1}(J_{i}/N)\right\}+\sum_{i=1}^{d+1}J_{i}\ln\left(\frac{s_{i}}{J_{i}/N}\right)+\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{J_{i}}$
@@
By the definition of $\widetilde{\gamma}(\boldsymbol{\varepsilon})$, one can write exactly
$\sum_{i=1}^{d+1}(J_{i}/N)\ln\left(\frac{p_{i}}{J_{i}/N}\right)=-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\widetilde{\gamma}(\boldsymbol{\varepsilon}).$ (10)
If, in addition, $\max_{i\in[d+1]}|\varepsilon_{i}|\leq\eta<1$, then the Taylor expansion
@@
and the fact that $\widetilde{\varepsilon}_{d+1}=-\sum_{i=1}^{d}\widetilde{\varepsilon}_{i}$ show that $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ can be expanded as follows:
@@
$=\frac{1}{2}\sum_{i,j=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\left\{\frac{1}{p_{i}}\mathds{1}\{i=j\}+\frac{1}{p_{d+1}}\right\}-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\frac{1}{6}\sum_{i,j,k=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\left\{\frac{1}{p_{i}^{2}}\mathds{1}\{i=j=k\}-\frac{1}{p_{d+1}^{2}}\right\}$
$\quad+\frac{1}{12}\sum_{i,j,k,\ell=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\widetilde{\varepsilon}_{\ell}\left\{\frac{1}{p_{i}^{3}}\mathds{1}\{i=j=k=\ell\}+\frac{1}{p_{d+1}^{3}}\right\}+\mathcal{O}_{d,\boldsymbol{p},\eta}\big(\|\boldsymbol{\varepsilon}\|_{1}^{5}\big).$
Moreover, by [15, Eq. 21], it is known that $(\Sigma_{\boldsymbol{p}}^{-1})_{ij}=p_{i}^{-1}\mathds{1}\{i=j\}+p_{d+1}^{-1}$ for all $i,j\in[d]$, so
$\frac{1}{2}\sum_{i,j=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\left\{\frac{1}{p_{i}}\mathds{1}\{i=j\}+\frac{1}{p_{d+1}}\right\}-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}=0.$ (11)
Under this additional condition, using the last two equations, one can rewrite $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ as follows:
$\widetilde{\gamma}(\boldsymbol{\varepsilon})$ $=-\frac{1}{6}\sum_{i,j,k=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\left\{\frac{1}{p_{i}^{2}}\mathds{1}\{i=j=k\}-\frac{1}{p_{d+1}^{2}}\right\}$ (12)
$\quad+\frac{1}{12}\sum_{i,j,k,\ell=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\widetilde{\varepsilon}_{\ell}\left\{\frac{1}{p_{i}^{3}}\mathds{1}\{i=j=k=\ell\}+\frac{1}{p_{d+1}^{3}}\right\}+\mathcal{O}_{d,\boldsymbol{p},\eta}\big(\|\boldsymbol{\varepsilon}\|_{1}^{5}\big).$
Now, consider $\boldsymbol{s}=(s_{1},\ldots,s_{d})\in\mathcal{R}_{d}^{\circ}$ as defined in (3). For all $i\in[d+1]$ and $t\in(0,1)$, let
$\delta_{i}(t)=\frac{t-p_{i}}{p_{i}}.$
In particular, note that this definition yields $\sum_{i=1}^{d+1}p_{i}\delta_{i}(s_{i})=0$. Upon applying Taylor’s formula with Lagrange remainder to the function $x\mapsto\ln(1+x)$, there exist points $s_{i}^{\star}$, with $s_{i}^{\star}$ between $s_{i}$ and $p_{i}$, $i\in[d+1]$, such that